# Mesh-Informed Neural Operator : A Transformer Generative Approach

### v1

@@ 2 Preliminaries
Table 1: Comparison of MINO with other architectures. A "local $\rightarrow$ global" receptive field, seen in models like U-Net, indicates that global information is gathered through a hierarchy of local operations. A detailed discussion is provided in Appendix C
@@
Figure 2: Inference and zero-shot Generation with MINO. (a) MINO gradually transforms a GP sample to the data sample under flow matching paradigm. (b) Zero-shot generation at varying spatial scales by directly transforms finer GP samples to finer data samples.
@@ 3 Methods
$\mathbf{f_{t}^{D}}=\text{concat}(\text{MLP}(\mathbf{f_{t}}),\mathbf{p_{emb}}))$ (6)
@@ 4 Experiments
Table 5: Computational efficiency and performance comparison between MINO variants and Transolver during training and inference.
@@ A.3 Hyperparameters for MINO
MINO-U. The GNO maps input functions to a latent representation on a $16\times 16$ grid of query points, defined over the $[0,1]^{2}$ domain. We set the GNO search radius to $0.07$, the latent dimension $L_{\text{dim}}$ to 256, and the number of attention heads to 4. The encoder consists of $M_{1}=2$ cross-attention blocks and the decoder has $M_{2}=2$ blocks. For its latent-space processor, we adopt a Diffusion U-Net architecture from the torchcfm library (Tong et al., 2024), which operates on the $[16,16]$ latent tensor with 64 channels, 1 residual block, and 8 attention heads for the processor. When applied to the Global Climate dataset, the number of U-Net channels is increased to 96 to accommodate the larger $32\times 16$ latent grid, while other settings remain the same. This configuration results in a total model size of 26.9 M for Global Climate dataset and 19.2 M for others.
@@ A.4 Details for training and inference
Reference Measures. We choose $\mu_{0}$ as a Gaussian measure characterized by a Gaussian Process (GP) with a Matérn kernel. Unless otherwise specified, the function domain is $[0,1]^{2}$ (or subset of it), and we use a kernel length scale of $0.01$ and a smoothness parameter of $0.5$. For the Global Climate dataset, the domain is a spherical manifold represented as a subset of $[-1,1]^{3}$, and the kernel length scale is adjusted to $0.05$. We utilize the GeometricKernels library (Mostowsky et al., 2024) for implementing the GP on the sphere. The variance is fixed at 1 for all GPs.
@@ C.2 Architectural Comparison to Other Neural Operators
Comparison with UPT. The UPT model uses a GNN for super-node pooling, a choice we identify as suboptimal for operator learning for two primary reasons. First, as the input discretization becomes finer, the GNN converges to a pointwise operator, which decouples it from the discretization convergence property required of a neural operator. Second, GNN can also be viewed as a low pass filter in the generative tasks, which has the same issue of GNO. Thrid, UPT relies on random subsampling to define super-nodes is ill-suited for adaptive meshes common in PDE tasks, where important regions with low grid density may be undersampled. For instance, on a dense mesh of the Earth, random subsampling would likely over-concentrate super-nodes in the polar regions that of less interest. Furthermore, UPT’s decoder uses a single-layer, perceiver-style cross-attention where the queries are derived only from the output positions. We found that for generative tasks, it is crucial to use a combination of the input function and its position as the query to the cross attention decoder and to use multiple attention layers as shown in the ablation study (Appendix E).

### v2

@@ 2 Preliminaries
Table 1: Comparison of MINO with other architectures. "local → global" receptive field gathers global context through stacked local operations. "local + global" receptive field combines local and global information directly. A detailed discussion is provided in Appendix C
@@
Figure 2: Inference and zero-shot Generation with MINO. (a) MINO gradually transforms a GP sample to the data sample under flow matching paradigm. (b) Zero-shot generation at varying spatial scales by directly transforming finer GP samples to finer data samples.
@@ 3 Methods
$\mathbf{f_{t}^{D}}=\text{MLP}(\text{concat}(\text{MLP}(\mathbf{f_{t}}),\mathbf{p_{emb}}))$ (6)
@@ 4 Experiments
Table 5: Computational efficiency and performance comparison between MINO variants and Transolver during training and inference on Shallow Water benchmark.
@@ A.3 Hyperparameters for MINO
MINO-U. The GNO maps input functions to a latent representation on a $16\times 16$ grid of query points, defined over the $[0,1]^{2}$ domain. We set the GNO search radius to $0.07$, the latent dimension $L_{\text{dim}}$ to 256, and the number of attention heads to 4. The encoder consists of $M_{1}=2$ cross-attention blocks and the decoder has $M_{2}=2$ blocks. For its latent-space processor, we adopt a Diffusion U-Net architecture from the torchcfm library (Tong et al., 2024), which operates on the $[16,16]$ latent tensor with 64 channels, 1 residual block, and 4 attention heads for the processor. When applied to the Global Climate dataset, the number of U-Net channels is increased to 96 to accommodate the larger $32\times 16$ latent grid, while other settings remain the same. This configuration results in a total model size of 26.9 M for Global Climate dataset and 19.2 M for others.
@@ A.4 Details for training and inference
Reference Measures. We choose $\mu_{0}$ as a Gaussian measure characterized by a Gaussian Process (GP) with a Matérn kernel. Unless otherwise specified, the function domain is $[0,1]^{2}$ (or subset of it), and we use a kernel length scale of $0.01$ and a smoothness parameter of $0.5$. For the Global Climate dataset, the domain is a spherical manifold represented as a subset of $[-1,1]^{3}$, and the kernel length scale is adjusted to $0.05$. Furthermore, for the GP on the sphere, we explored two distance metrics for the kernel. The first uses the chordal distance (the Euclidean distance in the $\mathbb{R}^{3}$ embedding space), while the second employs the geodesic distance on $\mathbb{S}^{2}$, for which we leverage the GeometricKernels library (Mostowsky et al., 2024). Although both approaches yield a valid GP, the results presented in this paper are based on the chordal distance implementation. We selected this method for two primary reasons: its superior computational speed and to ensure consistency with other benchmarks. To facilitate further exploration, we have made both implementations publicly available in our code repository. The variance is fixed at 1 for all GPs.
@@ C.2 Architectural Comparison to Other Neural Operators
Comparison with UPT. The UPT model uses a GNN for super-node pooling, a choice we identify as suboptimal for operator learning for three primary reasons. First, as the input discretization becomes finer, the GNN converges to a pointwise operator, which decouples it from the discretization convergence property required of a neural operator. Second, GNN can also be viewed as a low pass filter in the generative tasks, which has the same issue of GNO. Thrid, UPT relies on random subsampling to define super-nodes is ill-suited for adaptive meshes common in PDE tasks, where important regions with low grid density may be undersampled. For instance, on a dense mesh of the Earth, random subsampling would likely over-concentrate super-nodes in the polar regions that of less interest. Furthermore, UPT’s decoder uses a single-layer, perceiver-style cross-attention where the queries are derived only from the output positions. We found that for generative tasks, it is crucial to use a combination of the input function and its position as the query to the cross attention decoder and to use multiple attention layers as shown in the ablation study (Appendix E).